A dump truck hoists a box to an angle of 25° from the horizontal and pulls ahead 10 ft. Calculate the displacement of point A



1
Expert's answer
2022-04-12T02:48:03-0400

W = F × d × cos"\theta" = mg × d × cos"\theta" = mg × 10 ft × cos (25°)